The escape speed for a planet of mass M and radius R is 16.3 kms-1.

Calculate the escape velocity from a planet of mass four times the mass and twice the radius.


1
Expert's answer
2021-03-21T11:31:30-0400
"v=\sqrt{\frac{2GM}{r}}\\v'=\sqrt{\frac{2G(4)M}{2r}}=\sqrt{2}v\\v'=\sqrt{2}16.3=23.1\frac{km}{s}"